OpenStudy (anonymous):

Simplify. Write the answer in positive exponents. x^8/x^-9

OpenStudy (anonymous):

x^17

OpenStudy (anonymous):

\[x^8\div1/x ^{9}=x^8*x^9=x^{17}\]